A machine costs $40,000 to purchase and $10,000 per year to operate. The machine has no salvage life and a 10 year life. If i=10% per year compounded annually what is the present worth of the machine?

Find the present worth of the machine of Problem 7.8.

Ans. PW = -$I01 443.93 (the negative value indicates a cash outflow or cost)

Please show all work and formulas used thanks
Answer #1

Initial cost = 40000

Maintenance cost = 10000

n = 10

i = 10% = 0.1

The present worth

= -40000-10000*((1+i)^n-1)/(i(1+i)^n)

= -40000-10000*(1+0.1)^10-1)/0.1(1+0.1)^10)

= -40000-10000*(1.59374/)0.1(2.59374))

= -40000-10000*(1.59374/0.259374)

= -40000-10000*(6.14456)

= -40000-61445.6

= -101445.6